Session: Nov 1999 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Tue Nov 30 21:04:32 MET 1999) Minor failures
Some tracking problem during the observations due to antenna problems

Jodrell (Alastair Gunn at Wed Dec 1 03:41:51 MET 1999) Minor failures
Problem reading procedure file at start of experiment so first two scans not recorded properly. Telescope was parked at 14:38 UT due to high winds. Back on source at 19:00 UT. No other known problems.

Torun (K.Borkowski at Wed Dec 1 09:31:08 MET 1999) Success
No known problems.

Medicina ( at Wed Dec 1 14:58:45 MET 1999) Minor failures

Effelsberg (A. Kraus at Thu Dec 2 11:38:21 MET 1999) Success

Onsala (John Conway at Mon Dec 13 09:18:29 MET 1999) Minor failures
Only had useful data for approximately the first 45 minutes then had to stow due to winds.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:19:15 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: ep028.antab. Read the comments in the file to get further information.